OpenStudy (anonymous):

(2x+6)+130 = 180 solve for x to find angle

OpenStudy (dan815):

manipulate equation

OpenStudy (dan815):

go khanacademy.org he will teach you step by step

OpenStudy (anonymous):

(2x+6)+130=180 Simple and best practice solution for (2x+6)+130=180 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. Equation: Solution for (2x+6)+130=180 equation: Simplifying (2x + 6) + 130 = 180 Reorder the terms: (6 + 2x) + 130 = 180 Remove parenthesis around (6 + 2x) 6 + 2x + 130 = 180 Reorder the terms: 6 + 130 + 2x = 180 Combine like terms: 6 + 130 = 136 136 + 2x = 180 Solving 136 + 2x = 180 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-136' to each side of the equation. 136 + -136 + 2x = 180 + -136 Combine like terms: 136 + -136 = 0 0 + 2x = 180 + -136 2x = 180 + -136 Combine like terms: 180 + -136 = 44 2x = 44 Divide each side by '2'. x = 22

OpenStudy (anonymous):

thx bro

OpenStudy (anonymous):

Anytime..But ima girl..lol